Evening shift summary for 06/16/99

Patricia McBride (mcbride@fnpsph)
Thu, 17 Jun 1999 00:31:55 CDT

Shift Summary for Evening shift, 06/16/99

On Shift: Leo, Patty

Others present: RickF, Arun, Hogan, Sydney, Peter, and cast of thousands

Hours of beam delivered at all intensities: 8.0
Hours of beam delivered at nominal intensity: 8.0
Hours of E832 data (special runs not included): 3.0
Hours of special runs (laser scans, muon runs...): 1.5

Accounting of lost time:

Access to fix RC6 and to remove cable to DC/OR latches

========================================================================
Runs ended this shift:

Run 13627 - - BEAM_832
Run 13629 - Laser Scan -- aborted
Run 13630 - PED
Run 13631 - BEAM_832

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

17:30 Hogan, Arun and Leo make an access to fix DC6 power supply.
Could not raise shutter. Traced problem to a switch on the
door to KTeV Hall. After jiggling the door (not opening!), we were
able to raise the shutter.

21:00 Arun and Leo make another access to fix shaper splitter for RC6- CH1
Same problem with shutter. Same solution.

========================================================================
Overview of shift:

Goal(s) of shift: 1. E832 data-taking
2. Beam Studies by Rick Ford
3. Accesses to repair RC6


========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13627 15:30 Run in BEAM_832 mode in progress 4e12, BOA = 2-2.5

18:50 Stop run to investigate voltage problems on RC6.

JAG deletes 0 length RC files upto run 13625 for onsplit

Run control hangs while in process of stopping run.

do dart_fresh_start

19:00 Arun and Leo go to prep to get spare power supply
for RC6.
Rick Ford starts to tune the beam.

13628 17:30 Access to replace RC6 power supply.

Corrected a few problems with BA1 by moving
channels to different shpaer splitter module:

BA1 ch 10 --> BA2 ch 25
BA1 ch 74 --> BA2 ch 97
BA1 ch 77 --> BA2 ch 98

Hogan removed DCOR latch connections from DC1 and DC2 to
station 2.


20:26 Access complete, but RC6 channel 1 is not working. It
looks as if the problem is downstairs.

Rick Ford notices that NM2EU is not regulating properly.
MCR looking is into it.

20:57 Access to fix RC6 ch1. Installed new shaper splitter.
Moved RC6 ch1 over to the new module including trig.
sources.

21:45 Trouble ending run again. Do dart_fresh_start, logoff.
Do dart_fresh_start again.

13629 22:00 Start laser scan
22:10 Access over; abort laser scan.

13630 22:15 Beam is back; start pedestal run.

13631 22:20 RickF starts beam studies.

prolems remain with BA ch77; it might be the ADC.

22:50 Beam studies complete.
Results: 50 microrad easterly shift in tgt angle.

Call MCR to prepare muon run.



========================================================================
Items to Follow up on Next Shift

Call Elizabeth (3399) if there is a downtime of over 1.5 hrs.

Take a laser scan if beam is off for an hour or so.